OpenStudy (anonymous):

There are five seniors in a class. For each situation, write how the binomial formula is used to calculate the probability. a) In how many ways can I choose one senior to represent the group? b) In how many ways can I choose two seniors to represent the group? c) In how many ways can I choose three seniors to represent the group? d) In how many ways can I choose four seniors to represent the group? e) In how many ways can I choose five seniors to represent the group?

OpenStudy (anonymous):

@jim_thompson5910

OpenStudy (anonymous):

I got this but idk... :(a) n = 5 (5 seniors) and r = 1 (1 senior selected) --> 5 ways (b) (5 * 4) / 2 = 10 ways (c) 10 ways (d) 5 C(5-4) = 5C(1) = 5 ways (e) 1/1 = 1 way

jimthompson5910 (jim_thompson5910):

In how many ways can I choose one senior to represent the group? There are 5 ways because there are 5 seniors to choose from Say you had seniors code named A, B, C, D, and E There are only 5 ways to choose one letter

OpenStudy (anonymous):

ok i got a as 5 too.

OpenStudy (anonymous):

b?

OpenStudy (anonymous):

10 10 5 1 my answers....

jimthompson5910 (jim_thompson5910):

In how many ways can I choose two seniors to represent the group? There are 5 ways to choose the first senior There are 4 ways to choose the next so there are 5*4 = 20 ways to choose 2 seniors but since order doesn't matter, you have to divide by 2 to eliminate double counting so you're correct in saying (5*4)/2 = 10 ways

OpenStudy (anonymous):

ok so b is 10

jimthompson5910 (jim_thompson5910):

In how many ways can I choose three seniors to represent the group? There are 5*4*3 = 60 ways to choose 3 seniors But we overcount because order doesn't matter (ex: ABC is the same as CBA) so because there are 3! = 6 ways to order ABC, this means we must divide by 6 to get 60/6 = 10

OpenStudy (anonymous):

ok! so 10.. i got that too! now d?

jimthompson5910 (jim_thompson5910):

There are 5*4*3*2 = 120 different ways to pick 4 people (out of 5) Divide this by 4! = 24 to get 120/24 = 5 So that checks out as well

jimthompson5910 (jim_thompson5910):

Finally, there is only one way to make a group of 5 people when you only have 5 people to choose from (assuming order doesn't matter) So 5 C 5 = 1 In general, n C n = 1 for any positive integer n